OpenStudy (anonymous):

CAN YOU HELP??!! Find the value of the expression the square root of 6 squared + 8 squared - 5 squared. -15 -25 35 75

OpenStudy (anonymous):

Do you know how to figure out 6^2, 8 ^2, and 5^2?

OpenStudy (anonymous):

\[6 \times 6 + 8 \times 8 - 5 \times 5 =\]

OpenStudy (anonymous):

\(6 \times 6\) is bigger than \(5 \times 5\) so the whole thing must be bigger than \(8 \times 8\) so the anser bigger than that must be the right one